# Differentially Private Data Generation with Missing Data

### v2

@@ 1. Introduction
Our world as we see it today revolves around private data regarding our medical, financial, and social information. It is sometimes imperative to query such data for research and advancement of science (Kass et al., 2003; Andreou et al., 2013). Many industries also use statistics from private data to improve their products and user experience. However, reckless data sharing for data-driven applications and research causes great privacy concerns (Auxier et al., 2019; IBM, 2020) and penalties (GDP, 4 27). As a response, differential privacy (DP) (Dwork, 2006) has emerged as a standard data privacy guarantee which has now been adopted by government agencies (Abowd, 2018; Hawes, 2020) and companies (Erlingsson et al., 2014; Greenberg, 2016; Johnson et al., 2018). Informally, DP guarantees that the output distribution of an algorithm is similar with or without a particular individual in the dataset. A privacy budget is set to limit the total privacy loss and each query (e.g., releasing statistics (Abowd, 2018; Bureau, 1 30), building prediction models (Papernot et al., 2018; Abadi et al., 2016), and answering SQL queries (Ge et al., 2019; Johnson et al., 2018; Kotsogiannis et al., 2019; McSherry, 2009)) consumes part of the privacy budget, and once that budget is used up, no more queries can be answered directly. An alternative way is to generate a synthetic dataset using the privacy budget. The synthetic dataset, once generated, can be made public, and the analyst can use it for any number of downstream tasks  (Chen et al., 2018, 2015; Jordon et al., 2019; Zhang et al., 2014; Barak et al., 2007).
@@
Despite a number of work (Zhang et al., 2014; Ge et al., 2021; Phan et al., 2016; Jordon et al., 2019) that succeed in generating synthetic data with DP guarantees, they only look at a simple scenario where the input data has no missing values. Several prior studies (Little and Rubin, 2019; Getz et al., 2023) have reported on the prevalence of missing data in various fields. For instance, a study of 9 publicly available healthcare datasets commonly used in machine learning research found that the proportion of missing values ranged from 0.2% to 78.6% (Getz et al., 2023). The presence of missing data can be attributed to multiple reasons, such as human errors (Agarwal and Singh, 2021) and privacy regulations like GDPR (GDP, 4 27) which allow people the “right to forget” where one may ask their data to be deleted completely (Politou et al., 2018). In our work, we ask how missing data will affect the quality of the synthetic data generated by mechanisms that offer DP guarantees. Our preliminary study shows that existing DP mechanisms have 4%-18.5% decay in the F1-score of downstream ML tasks on the synthetic dataset generated from a dataset with 10% missing values as compared to that when generated from the complete dataset. The decay varies depending on the types of missing mechanisms and the types of data generation processes.
@@
In our work, we formally define the research problem of generating synthetic data for sensitive data with missing values using DP. We consider a missing mechanism that takes complete ground truth data and outputs data with missing values. Under this setup, we can offer DP to either the incomplete data or the ground truth data. For each privacy guarantee, we study how to handle missing data in the synthetic data generation process. While it is desired to have a unified comprehensive approach to solve the problem, we identify multiple challenges in achieving such a solution. First, there is no straightforward winning algorithm for DP synthetic data generation even without missing data as demonstrated by prior benchmarking work (Tao et al., 2021). Second, the design space for a solution is huge for dealing with missing data. We explore several techniques, including the vanilla approach that uses complete rows only (Nakagawa and Freckleton, 2008), common imputation techniques (e.g., statistical methods, machine learning methods) (Little and Rubin, 2019; Rubin, 1976; Nordholt, 1998; Lakshminarayan et al., 1996), and other differentially private imputation methods (Das et al., 2022; Ge et al., 2021). We show that all these methods have their limitations, such as discarding too many rows or incurring high privacy costs, leading to poor-quality synthetic data. Therefore in our work we present a comprehensive list of possible approaches to solve the problem including two vanilla approaches and three adaptive approaches, as a contribution to understanding the design space for this new problem.
@@
In addition, we show the relationship between the privacy guarantee for these DP mechanisms that they offer for the incomplete data and that for the ground truth data. To do so, we model the missing mechanism as a sampling process and obtain a tighter upper bound for the privacy loss to the ground truth data via sampling amplification techniques (Balle et al., 2018; Steinke, 2022). Unlike prior work for sampling amplification that considers a random subset, we make use of the randomness due to missing values to amplify the privacy for ground truth data. The major contributions of our work are as follows:
@@ 2.1. Missing Data
Missing data is classified into different types using missing mechanisms. A missing mechanism $M_{\phi}:\mathcal{D}\rightarrow\mathcal{D}$ takes as input the ground truth dataset $\bar{D}$ and outputs an incomplete dataset $D$. It is parameterized by $\Phi$, a set of probabilities, which refers to the set of probabilities that control the unknown missing data process. Three missing types can be defined using $\Phi$ and the conditional distribution of missing indicator $m_{i}$ given the dataset $D_{i}$  (Little and Rubin, 2019; Rubin, 1976).
Missing completely at random (MCAR) assumes the probability of missingness is completely independent of the data. Under MCAR, any two rows of the dataset, regardless of their values, for the same attribute have the same probability of having a missing value. Hence, the parameter set $\Phi$ consists of $\{\phi_{j}|j\in[1,k]$}, where $\phi_{j}$ is the probability of any row having a missing value for the $j$th column. For $j\in[1,k]$ and $i\in[1,n]$, $\Pr[m_{ij}|D]=\Pr[m_{ij}]=\phi_{j}$. Hence, the probability of a row having no missing values is $\prod_{j=1}^{k}(1-\phi_{j})$.
@@
Datasets often contain various missing data types, but identifying them without domain knowledge remains challenging (Little, 1988, 1994). While statistical tests exist for MCAR cases (e.g., Little’s MCAR test, pattern mixture models) (Little, 1988, 1994), identifying MAR and MNAR remains unsolved due to complex interactions between observed and unobserved variables (Van Buuren, 2018). Despite binary indicator tests based on modeling for MAR, these lack conclusiveness and are sensitive to analysis assumptions. For private datasets, these tests also need to be computed privately by using a portion of the privacy budget.
@@ 2.2. Differential Privacy
Differential privacy (DP) (Dwork et al., 2006b, a) is used as our measure of privacy.
Definition 2.1 (Differential Privacy (DP) (Dwork and Roth, 2014; Dwork et al., 2006a)).
A randomized algorithm $M$ achieves ($\epsilon,\delta$)-DP if for all $Z\subseteq$ $\mathcal{M}$) and for two neighboring databases $D,D^{\prime}\in\mathcal{D}$ that differ in one row:
@@
Gaussian mechanism (Dwork and Roth, 2014) and Laplace mechanism (Dwork et al., 2006b) are two such widely used DP algorithms. Given a function $f:\mathcal{D}\rightarrow\mathbb{R}^{d}$, the Gaussian mechanism adds noise sampled from a Gaussian distribution $\mathcal{N}(0,S_{f}^{2}\sigma^{2})$ to each component of the query output, where $\sigma$ is the noise scale and $S_{f}$ is the $L_{2}$ sensitivity of function $f$, which is defined as $S_{f}=\max_{D,D^{\prime}\text{differ in a row}}||f(D)-f(D^{\prime})||_{2}$. For $\epsilon\in(0,1)$, if $\sigma\geq\sqrt{2\ln(1.25/\delta)}/\epsilon$, the Gaussian mechanism satisfies $(\epsilon,\delta)$-DP. Laplace mechanism works similarly but with the noise from the Laplace distribution and the $L_{1}$ sensitivity. Both these mechanisms have been applied to answer counting queries (Li et al., 2015) and is widely used in estimating low dimensional statistics about the dataset.
Complex DP algorithms can be built from these basic algorithms following two important properties of DP: 1) Post-processing (Dwork et al., 2006a) states that for any function $g$ defined over the output of the mechanism $\mathcal{M}$, if $\mathcal{M}$ satisfies ($\epsilon,\delta$)-DP, so does $g(\mathcal{M})$; 2) Composability (Dwork, 2006) states that if $\mathcal{M}_{1}$, $\mathcal{M}_{2}$, $\cdots$, $\mathcal{M}_{k}$ satisfy ($\epsilon_{1},\delta_{1}$)-, ($\epsilon_{1},\delta_{1}$)-, $\cdots$, ($\epsilon_{k},\delta_{k}$)-DP, then sequentially applying these mechanisms satisfies ($\sum_{i=1}^{k}\epsilon_{i},\sum_{i=1}^{k}\delta_{i}$)-DP.
@@ 2.3. DP Synthetic Data Generation
A common DP study is to generate synthetic data given a fixed privacy budget. The synthetic data, once generated, can be made public and all queries on this dataset come for free due to the post-processing property of DP. There are three main approaches for DP synthetic data generation (Tao et al., 2021):
Statistical approaches rely on estimating low-dimensional statistics about the dataset such as marginals (Xiao et al., 2011; Qardaji et al., 2014; McKenna et al., 2022). These approaches can be made better by finding the correlation between attributes. Techniques for improvement include probabilistic models (Koller and Friedman, 2009), Bayesian models (Zhang et al., 2014; Li et al., 2014; Ping et al., 2017) and undirected graphs (Chen et al., 2015; McKenna et al., 2019). Statistical approaches capture the underlying distribution of the correlated independent attributes very well but fail to imbibe complex relationships between multiple attributes.
Deep learning approaches are promising for generating synthetic data (Shrivastava et al., 2017; Chawla, 2019; Gupta, 2019), particularly with autoencoders and generative adversarial networks (GAN). Autoencoders map data into a low-dimensional feature space and sample synthetic data from the low-dimensional space. GANs utilize a generator to produce fake examples and a discriminator to distinguish real from fake. DPSGD (Williams and McSherry, 2010; Bassily et al., 2014; Song et al., 2013; Abadi et al., 2016) is commonly used for privacy. Various private approaches for autoencoders (Phan et al., 2016; Abay et al., 2018; Acs et al., 2018) and GANs (Frigerio et al., 2019; Jordon et al., 2019; Torkzadehmahani et al., 2020; Xie et al., 2018) have been proposed, effective on image data but challenged with tabular data due to poor encoding. Conditional GANs (Xu et al., 2019) and private versions (Torkzadehmahani et al., 2019) address this by sampling based on conditional probabilities of categorical attributes.
Mixed approaches are inspired by both the above approaches and try to preserve both low-dimensional statistics and high-level information. Some techniques include leveraging the dimensionality reduction via random orthonormal (RON) projection, the Gaussian generative model (Chanyaswad et al., 2019), combining denial constraints and attribute-wise embedding models (Ge et al., 2021) and Gretel.ai statistics (Noruzman et al., [n.d.]).
@@ 3. Problem Statement
In Section 4, we delineate multiple options for the data curator to deal with incomplete data, put forward challenges that come with them, and discuss which option might be the best and when. The first option is a naïve adaptation of prior work for DP data generation by simply discarding the rows with missing values (Rubin, 1976). We refer to this approach as complete row only. This approach can fail in many cases, as detailed in Section 4.1. For instance, if all rows have some missing values, then there will be no input data for the data generation methods. A second approach is to impute the missing parts with inferred values from the observed data. We denote this approach as imputation first approach. However, as our data is private, the imputation process needs to be privatized as well and the additional incurred privacy cost must be accounted for in the privacy budget. In section 4.1, we explore the privacy costs of imputation and show how they can be expensive in practice. Rather than having separate processes for imputation and synthetic data generation, we can integrate these two processes into one. This line of thought motivates us to a new approach, which we call adaptive recourse approach. In Section 4.2, we improve upon three categories of DP generation approaches and demonstrate their effectiveness in generating synthetic data from incomplete data. These strategies use no extra privacy budget and solely improve by observing available data in the dataset.
@@
[Privacy for Ground Truth Data] Consider the same setup as Problem 1. The trusted data curator uses the incomplete dataset $D$ as input and aims to generate a synthetic data $D^{*}$ of $n$ rows with a mechanism $M:\mathcal{D}\rightarrow\mathcal{D}$ such that minimizes $d(f(\bar{D})-f(D^{*}))$, where $f:\mathcal{D}\rightarrow\mathbb{R}^{l}$ is any utility metric function that the user is interested in, $d(\cdot,\cdot)$ is a distance metric and $M\circ M_{\Phi}$ offers $(\bar{\epsilon},\bar{\delta})$-DP to the ground truth data $\bar{D}$.
@@ 4.1. Vanilla Approaches
    Statistical methods: Each attribute of the dataset is modeled separately using statistical methods such as mean, median, and mode. The model is then used to fill up the missing values of the attribute (Little and Rubin, 2019; Rubin, 1976). For example, in Figure 2, for the left tables, we use the median of the observed values of column A, to fill up the missing value of the 4th row.
@@
    Hot and cold deck methods: This imputation technique replaces every missing value with another value from the same dataset (Hot deck) or from a proxy dataset (Cold deck) (Andridge and Little, 2010; Nordholt, 1998). The missing cells of the incomplete row are then filled up from the closest similar row. Similarity metric like cosine distance or $\ell_{2}$ distance can be used.
@@
    ML imputation: Machine learning (ML) based approaches are common for missing data imputation (Jerez et al., 2010; Lakshminarayan et al., 1996). An ML model is trained to predict the missing values of an attribute based on other non-missing attributes in the dataset as training features.
@@
(McSherry, 2009) We say a transformation $T(\cdot)$ is $c$-stable, if the distance between $T(D)$ and $T(D^{\prime})$ is at most $c$ times the distance between $D$ and $D^{\prime}$. The composite mechanism $\mathcal{M}\circ T$ then becomes $(c\cdot\epsilon,\delta)$-DP, for any mechanism $\mathcal{M}$ which is $(\epsilon,\delta)$-DP.

### v3

@@ 1. Introduction
Our world as we see it today revolves around private data regarding our medical, financial, and social information. It is sometimes imperative to query such data for research and advancement of science (kass2003use; andreou2013should). Many industries also use statistics from private data to improve their products and user experience. However, reckless data sharing for data-driven applications and research causes great privacy concerns (user_concern; ibm_report) and penalties (GDPR). As a response, differential privacy (DP) (DBLP:conf/icalp/Dwork06) has emerged as a standard data privacy guarantee which has now been adopted by government agencies (DBLP:conf/kdd/Abowd18; Hawes2020Implementing) and companies (Erlingsson14Rappor; Greenberg16Apples; DBLP:journals/pvldb/JohnsonNS18). Informally, DP guarantees that the output distribution of an algorithm is similar with or without a particular individual in the dataset. A privacy budget is set to limit the total privacy loss and each query (e.g., releasing statistics (DBLP:conf/kdd/Abowd18; onthemap), building prediction models (DBLP:conf/iclr/PapernotSMRTE18; DBLP:conf/ccs/AbadiCGMMT016), and answering SQL queries (DBLP:conf/sigmod/0002HIM19; DBLP:journals/pvldb/JohnsonNS18; DBLP:journals/pvldb/KotsogiannisTHF19; DBLP:conf/sigmod/McSherry09)) consumes part of the privacy budget, and once that budget is used up, no more queries can be answered directly. An alternative way is to generate a synthetic dataset using the privacy budget. The synthetic dataset, once generated, can be made public, and the analyst can use it for any number of downstream tasks  (DBLP:journals/corr/abs-1812-02274; DBLP:conf/kdd/ChenXZX15; DBLP:conf/iclr/JordonYS19a; DBLP:conf/sigmod/ZhangCPSX14; DBLP:conf/pods/BarakCDKMT07).
Despite a number of work (DBLP:conf/sigmod/ZhangCPSX14; ge2021kamino; phan2016differential; DBLP:conf/iclr/JordonYS19a) that succeed in generating synthetic data with DP guarantees, they only look at a simple scenario where the input data has no missing values. Several prior studies (little2019statistical; getz2023performance) have reported on the prevalence of missing data in various fields.
@@
For instance, a study of 9 publicly available healthcare datasets commonly used in machine learning research found that the proportion of missing values ranged from 0.2% to 78.6% (getz2023performance). The presence of missing data can be attributed to multiple reasons, such as human errors (agarwal2021causal) and privacy regulations like GDPR (GDPR) which allow people the “right to forget” where one may ask their data to be deleted completely (politou2018forgetting).
In our work, we ask how missing data will affect the quality of the synthetic data generated by mechanisms that offer DP guarantees. Our preliminary study shows that existing DP mechanisms have 4%-18.5% decay in the F1-score of downstream ML tasks on the synthetic dataset generated from a dataset with 10% missing values as compared to that when generated from the complete dataset. The decay varies depending on the types of missing mechanisms and the types of data generation processes.
@@
In our work, we formally define the research problem of generating synthetic data for sensitive data with missing values using DP. We consider a missing mechanism that takes complete ground truth data and outputs data with missing values. Under this setup, we can offer DP to either the incomplete data or the ground truth data. For each privacy guarantee, we study how to handle missing data in the synthetic data generation process. While it is desired to have a unified comprehensive approach to solve the problem, we identify multiple challenges in achieving such a solution. First, there is no straightforward winning algorithm for DP synthetic data generation even without missing data as demonstrated by prior benchmarking work (tao2021benchmarking). Second, the design space for a solution is huge for dealing with missing data. We explore several techniques, including the vanilla approach that uses complete rows only (nakagawa2008missing), common imputation techniques (e.g., statistical methods, machine learning methods) (little2019statistical; rubin1976inference; nordholt1998imputation; lakshminarayan1996imputation), and other differentially private imputation methods (DBLP:journals/corr/abs-2206-15063; ge2021kamino). We show that all these methods have their limitations, such as discarding too many rows or incurring high privacy costs, leading to poor-quality synthetic data. Therefore in our work we present a comprehensive list of possible approaches to solve the problem including two vanilla approaches and three adaptive approaches, as a contribution to understanding the design space for this new problem.
@@
In addition, we show the relationship between the privacy guarantee for these DP mechanisms that they offer for the incomplete data and that for the ground truth data. To do so, we model the missing mechanism as a sampling process and obtain a tighter upper bound for the privacy loss to the ground truth data via sampling amplification techniques (balle2018privacy; steinke2022composition). Unlike prior work for sampling amplification that considers a random subset, we make use of the randomness due to missing values to amplify the privacy for ground truth data. The major contributions of our work are as follows:
@@ 2.1. Missing Data
Missing data is classified into different types using missing mechanisms. A missing mechanism $M_{\phi}:\mathcal{D}\rightarrow\mathcal{D}$ takes as input the ground truth dataset $\bar{D}$ and outputs an incomplete dataset $D$. It is parameterized by $\Phi$, a set of probabilities, which refers to the set of probabilities that control the unknown missing data process. Three missing types can be defined using $\Phi$ and the conditional distribution of missing indicator $m_{i}$ given the dataset $D_{i}$  (little2019statistical; rubin1976inference).
Missing completely at random (MCAR) assumes the probability of missingness is completely independent of the data. Under MCAR, any two rows of the dataset, regardless of their values, for the same attribute have the same probability of having a missing value. Hence, the parameter set $\Phi$ consists of $\{\phi_{j}|j\in[1,k]$}, where $\phi_{j}$ is the probability of any row having a missing value for the $j$th column.
For $j\in[1,k]$ and $i\in[1,n]$, $\Pr[m_{ij}|D]=\Pr[m_{ij}]=\phi_{j}$.
Hence, the probability of a row having no missing values is $\prod_{j=1}^{k}(1-\phi_{j})$.
@@
Datasets often contain various missing data types, but identifying them without domain knowledge remains challenging (little1988test; little1994class). While statistical tests exist for MCAR cases (e.g., Little’s MCAR test, pattern mixture models) (little1988test; little1994class), identifying MAR and MNAR remains unsolved due to complex interactions between observed and unobserved variables (van2018flexible). Despite binary indicator tests based on modeling for MAR, these lack conclusiveness and are sensitive to analysis assumptions. For private datasets, these tests also need to be computed privately by using a portion of the privacy budget.
@@ 2.2. Differential Privacy
Differential privacy (DP) (DworkMNS06; DBLP:conf/eurocrypt/DworkKMMN06) is used as our measure of privacy.
Definition 2.1 (Differential Privacy (DP) (DBLP:journals/fttcs/DworkR14; DBLP:conf/eurocrypt/DworkKMMN06)).
A randomized algorithm $M$ achieves ($\epsilon,\delta$)-DP if for all $Z\subseteq$ Range($\mathcal{M}$) and for two neighboring databases $D,D^{\prime}\in\mathcal{D}$ that differ in one row:
@@
[…]
Complex DP algorithms can be built from these basic algorithms following two important properties of DP: 1) Post-processing (DBLP:conf/eurocrypt/DworkKMMN06) states that for any function $g$ defined over the output of the mechanism $\mathcal{M}$, if $\mathcal{M}$ satisfies ($\epsilon,\delta$)-DP, so does $g(\mathcal{M})$; 2) Composability (DBLP:conf/icalp/Dwork06) states that if $\mathcal{M}_{1}$, $\mathcal{M}_{2}$, $\cdots$, $\mathcal{M}_{k}$ satisfy ($\epsilon_{1},\delta_{1}$)-, ($\epsilon_{1},\delta_{1}$)-, $\cdots$, ($\epsilon_{k},\delta_{k}$)-DP, then sequentially applying these mechanisms satisfies ($\sum_{i=1}^{k}\epsilon_{i},\sum_{i=1}^{k}\delta_{i}$)-DP.
@@ 2.3. DP Synthetic Data Generation
A common DP study is to generate synthetic data given a fixed privacy budget. The synthetic data, once generated, can be made public and all queries on this dataset come for free due to the post-processing property of DP. There are three main approaches for DP synthetic data generation (tao2021benchmarking):
Statistical approaches rely on estimating low-dimensional statistics about the dataset such as marginals (DBLP:journals/tkde/XiaoWG11; DBLP:conf/sigmod/QardajiYL14; mckenna2022aim). These approaches can be made better by finding the correlation between attributes. Techniques for improvement include probabilistic models (DBLP:books/daglib/0023091), Bayesian models (DBLP:conf/sigmod/ZhangCPSX14; DBLP:journals/pvldb/LiXZJ14; DBLP:conf/ssdbm/PingSH17) and undirected graphs (DBLP:conf/kdd/ChenXZX15; DBLP:conf/icml/McKennaSM19). Statistical approaches capture the underlying distribution of the correlated independent attributes very well but fail to imbibe complex relationships between multiple attributes.
Deep learning approaches are promising for generating synthetic data (DBLP:conf/cvpr/ShrivastavaPTSW17; Chawla2019DeepfakesH; DBLP:conf/icassp/Gupta19), particularly with autoencoders and generative adversarial networks (GAN). Autoencoders map data into a low-dimensional feature space and sample synthetic data from the low-dimensional space. GANs utilize a generator to produce fake examples and a discriminator to distinguish real from fake. DPSGD (WM10; BST14; song2013stochastic; DBLP:conf/ccs/AbadiCGMMT016) is commonly used for privacy. Various private approaches for autoencoders (phan2016differential; abay2018privacy; acs2018differentially) and GANs (DBLP:conf/sec/FrigerioOGD19; DBLP:conf/iclr/JordonYS19a; DBLP:journals/corr/abs-2001-09700; DBLP:journals/corr/abs-1802-06739) have been proposed, effective on image data but challenged with tabular data due to poor encoding. Conditional GANs (xu2019modeling) and private versions (torkzadehmahani2019dp) address this by sampling based on conditional probabilities of categorical attributes.
Mixed approaches are inspired by both the above approaches and try to preserve both low-dimensional statistics and high-level information. Some techniques include leveraging the dimensionality reduction via random orthonormal (RON) projection, the Gaussian generative model (chanyaswad2019ron), combining denial constraints and attribute-wise embedding models (ge2021kamino) and Gretel.ai statistics (noruzmangretel).
@@ 3. Problem Statement
In Section 4, we delineate multiple options for the data curator to deal with incomplete data, put forward challenges that come with them, and discuss which option might be the best and when. The first option is a naïve adaptation of prior work for DP data generation by simply discarding the rows with missing values (rubin1976inference). We refer to this approach as complete row only. This approach can fail in many cases, as detailed in Section 4.1. For instance, if all rows have some missing values, then there will be no input data for the data generation methods. A second approach is to impute the missing parts with inferred values from the observed data. We denote this approach as imputation first approach. However, as our data is private, the imputation process needs to be privatized as well and the additional incurred privacy cost must be accounted for in the privacy budget. In section 4.1, we explore the privacy costs of imputation and show how they can be expensive in practice. Rather than having separate processes for imputation and synthetic data generation, we can integrate these two processes into one. This line of thought motivates us to a new approach, which we call adaptive recourse approach. In Section 4.2, we improve upon three categories of DP generation approaches and demonstrate their effectiveness in generating synthetic data from incomplete data. These strategies use no extra privacy budget and solely improve by observing available data in the dataset.
@@
[Privacy for Ground Truth Data] Consider the same setup as Problem 1.
The trusted data curator uses the incomplete dataset $D$ as input and aims to generate a synthetic data $D^{*}$ of $n$ rows with a mechanism $M:\mathcal{D}\rightarrow\mathcal{D}$ such that minimizes $d(f(\bar{D})-f(D^{*}))$, where $f:\mathcal{D}\rightarrow\mathbb{R}^{l}$ is any utility metric function that the user is interested in, $d(\cdot,\cdot)$ is a distance metric and $M\circ M_{\Phi}$ offers $(\bar{\epsilon},\bar{\delta})$-DP to the ground truth data $\bar{D}$.
@@ 4.1. Vanilla Approaches
    Statistical methods: Each attribute of the dataset is modeled separately using statistical methods such as mean, median, and mode. The model is then used to fill up the missing values of the attribute (little2019statistical; rubin1976inference). For example, in Figure 2, for the left tables, we use the median of the observed values of column A, to fill up the missing value of the 4th row.
@@
    Hot and cold deck methods: This imputation technique replaces every missing value with another value from the same dataset (Hot deck) or from a proxy dataset (Cold deck) (andridge2010review; nordholt1998imputation).
    The missing cells of the incomplete row are then filled up from the closest similar row. Similarity metric like cosine distance or $\ell_{2}$ distance can be used.
@@
    ML imputation: Machine learning (ML) based approaches are common for missing data imputation (jerez2010missing; lakshminarayan1996imputation). An ML model is trained to predict the missing values of an attribute based on other non-missing attributes in the dataset as training features.
@@
(DBLP:conf/sigmod/McSherry09) We say a transformation $T(\cdot)$ is $c$-stable, if the distance between $T(D)$ and $T(D^{\prime})$ is at most $c$ times the distance between $D$ and $D^{\prime}$. The composite mechanism $\mathcal{M}\circ T$ then becomes $(c\cdot\epsilon,\delta)$-DP, for any mechanism $\mathcal{M}$ which is $(\epsilon,\delta)$-DP.